# disposition

### v2

@@ 3 Desirable Properties of Dependence Measures
Usually, directed dependence measures lie in the interval $[-1,1]$ (property (A)), where negative values indicate negative and positive values indicate positive dependence. (B) and (C) make sure that $\delta(A,B)$ correctly indicates the extreme cases of independence as well as perfect positive and negative dependence. Note that, contrary to the case of two general random variables, a dependence measure in the binary case is not a summary measure in the sense of having to condense information (a complicated dependence structure characterized fully by a joint CDF or copula) into a single number, but it usually characterizes the full dependence structure in this single number in the sense that given the marginal event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$ and the dependence measure $\delta(A,B)$, we can recover $\mathbb{P}(A\cap B)$ (indeed, this is possible for all measures discussed in this paper except for the contingency coefficients). As a consequence of this, the “only if” in (B) is infeasible for directed measures of dependence for arbitrary random variables (Embrechts et al.,, 2002), but in the binary case it is feasible. (D) is a very natural property, making sure that stronger dependence leads to more extreme values of $\delta(A,B)$. It is related to the property of coherence proposed by Scarsini, (1984). (C) and (D) ensure that a dependence measure is a sensible indicator of strength of dependence in that larger absolute values of the measure indicate stronger dependence and absolute values close to 0 and 1 indicate weak and strong dependence, respectively. Properties (B) and (D) also imply that the measure correctly indicates the direction of dependence as formalized through the following Proposition.
@@
[…]
@@ 4.1 Covariance
They imply $-\frac{1}{4}\leq{Cov}(A,B)\leq\frac{1}{4}$. Thus, property (A) is fulfilled, but property (C) is not. Properties (B) and (D) follow by definition of independence and stronger positive (negative) dependence, (E) directly by representation (1).
@@
${Cov}(A,B)$ fulfills properties (A), (B), (D) and (E) from Definition 3.2, but not (C).
Thus, if one found a suitable normalization that made the covariance attainable, a proper dependence measure would arise. Indeed, the phi coefficient, Cole’s ${C}$ and Yule’s ${Q}$, which we discuss below, all turn out to be normalized versions of the covariance.
@@ 4.2 Phi Coefficient
The phi coefficient is just the classical Pearson correlation coefficient (as well as Spearman’s $\rho$ and Kendall’s $\tau_{b}$, which all coincide in this setting) of $\mathds{1}_{A}$ and $\mathds{1}_{B}$ and thus a normalized version of ${Cov}(A,B)$.
@@
As these bounds are in general not equal to $-1$ and $1$ and depend on the marginal event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$, the phi coefficient is not attainable.
@@
[…]
@@
Thus, the phi coefficient is not a particularly useful measure of dependence for events or binary random variables. Instead, it may be useful for a different task, which becomes clear by analyzing under which conditions $\phi$ takes the values $-1$ and 1.
@@
    $A$ and $B$ are perfectly positively (negatively) dependent and $\mathbb{P}(A)=\mathbb{P}(B)$ ($\mathbb{P}(A)=\mathbb{P}(\overline{B})$).
@@
What has been recognized in the literature (e.g. Bishop et al., (2007)), is that the phi coefficient equals 1($-1$) under condition (iv), that is, if there are two zeros on the primary (secondary) diagonal of the contingency table, while other measures like Yule’s ${Q}$ (in fact, all proper measures by (C) from Definition 3.2 and (iii) in Proposition 2.4) require only one zero on the primary (secondary) diagonal to equal 1 (--1). However, the underlying reasons for this as made clear by Propositions 2.4 and 4.5 have not been uncovered.
The items (vi) and (vii) in Proposition 4.5 are interesting because they establish the connection to conditions well-known from the theory of Pearson correlation and rank correlations, of which $\phi$ is a special case. (vii) reproduces the well-known condition that Pearson correlation is 1 (--1) if and only if the two variables are linear functions of each other, (vi) the well-known condition that Spearman’s $\rho$ and Kendall’s $\tau_{b}$ are 1 (--1) if and only if one variable is a monotonic function of the other (Embrechts et al.,, 2002). Thus, equality between the events (or one event and the complement of the other) corresponds to positive (negative) monotonic and linear (the two coincide in the binary case) dependence. Note that neither linear nor monotonic functional relations are useful concepts of perfect dependence in the binary case (or more generally in the discrete case) as a functional relation is only possible under the restriction from (iii) on the marginal distributions, which is in general not fulfilled.
@@ 4.3 Contingency Coefficients
Contingency coefficients aim at measuring dependence inherent in a general $k\times l$ contingency table. For two categorical random variables $X$ and $Y$, where $X$ takes values $x_{1},...,x_{k}$ and $Y$ takes values $y_{1},...,y_{l}$, Pearson’s Mean Square Contingency (${MSC}$) coefficient (Pearson,, 1904) is just the population analogue of the test statistic of Pearson’s chi-squared test of independence:
@@
Contingency coefficients arise by normalizing ${MSC}(X,Y)$. As they are designed for nominal random variables, direction of dependence is not a sensible concept and they map to $[0,1]$, only trying to indicate strength of dependence. However, as discussed above, binary random variables can always be regarded as ordinal. Thus, in the $2\times 2$ case, measures of dependence that indicate strength and direction of dependence can be used and the question arises why one should throw away the information on direction of dependence. Since they are widely-used in the binary case, we nevertheless discuss them. In Appendix A.1, we consider Cramérs V, Tschuprow’s T, and Pearson’s contingency coefficient and show that they are closely related to the phi coefficient in the binary setting (for example the first two just equal $|\phi(A,B)|$) and thus inherit its deficiencies, in particular not being proper dependence measures.
@@ 5.1 Cole’s Coefficient
Cole’s Coefficient was proposed by Cole, (1949). An equivalent reformulation of it arises naturally from our discussion of dependence concepts: From the Fréchet–Hoeffding bounds of $\mathbb{P}(A\cap B)$ from Proposition 2.4 arise the bounds on ${Cov}(A,B)$ from (2), which are attained only under perfect positive and negative dependence of $A$ and $B$ and whose absolute values are in general different from each other. Thus, it is natural to distinguish the cases of positive and negative dependence and normalize covariance with (the absolute values of) those bounds.
@@
The idea of normalizing with the cases of perfect positive and negative dependence has later been used for rank correlations (Vandenhende and Lambert, (2003), Genest and Nešlehová, (2007)) and for Pearson correlation and generalizations of it by Fissler and Pohle, (2023). Indeed, Cole’s ${C}$ is a special case of attainable modifications of Kendall’s tau, Spearman’s rho and Pearson correlation as well as of threshold correlation (see Genest and Nešlehová, (2007) for the former two and Fissler and Pohle, (2023) for the latter two).
@@ 5.2 Yule’s Q, the Odds Ratio and Relatives
Besides the phi coefficient and Cole’s coefficient, Yule’s ${Q}$ can be viewed as yet another approach to normalizing covariance. It just replaces the minus in the alternative representation of ${Cov}(A,B)$ from (1) with a plus and normalizes by the resulting term.
Definition 5.3 (Yule’s ${Q}$).
Yule’s ${Q}$ is defined as
@@
Considering (6), it is clear that ${Q}(A,B)$ lies in $[-1,1]$ and fulfills the symmetry properties (E). Its attainability directly follows by invoking part (iii) of Proposition 2.4. The independence property (B) is inherited from the covariance (since the normalization is nonzero). Monotonicity follows from the strictly monotonic relation between Yule’s ${Q}$ and the odds ratio (Lemma 5.6) and the monotoniticy of the odds ratio (part (D) of Proposition A.3).
@@
${Q}(A,B)$ is proper.
The normalization of Yule’s ${Q}$ is rather ad hoc, in particular compared to the natural normalization of Cole’s ${C}$. The normalization does not use the Fréchet–Hoeffding bounds for covariance, but continuously changes with the joint probabilities. Thus, the measure is not as nicely interpretable in terms of covariance relative to its values under perfect positive or negative dependence. On the other hand, it avoids the need to use a case distinction in the normalization as Cole’s ${C}$ does, which makes asymptotic theory and inference much easier (see Section 6). One way to interpret Yule’s ${Q}$ is by realizing that it is a special case of Goodman-Kruskal’s $\gamma$ (Goodman and Kruskal,, 1954) and can therefore be interpreted in the same vein (Bishop et al.,, 2007).
A very natural way to motivate and interpret Yule’s ${Q}$ is via its relation to the odds ratio. The odds ratio is different from the other measures of dependence discussed so far in that it does not build on the covariance and does not fall into the class of classical dependence measures mapping to $[-1,1]$ or $[0,1]$ as it maps to the extended nonnegative real numbers. Even though it thus falls a bit out of the framework of this paper, we nevertheless discuss it here shortly due to its popularity and to it being nicely interpretable and having nice properties.
@@
We discuss further details on the motivation behind the odds ratio, its interpretation and popularity in Appendix A.2. Note that the odds ratio is closely related to the covariance in that it just replaces the subtraction in (1) by a division. Fittingly, it is also sometimes called the cross-product ratio, referring to the underlying contingency table. The odds ratio and Yule’s ${Q}$ are closely related: They are strictly monotonic functions of each other.
@@
Thus, Yule’s ${Q}$ can be viewed as a transformation of the odds ratio to a new scale. Hence, it is unsurprising that the odds ratio is proper in an appropriate sense as well. Indeed, we show that it fulfills a modified set of axioms adopted to its codomain and its multiplicative rather than additive nature in Proposition A.3 in the appendix.
Furthermore, Yule’s ${Q}$ has been generalized in the sense of
@@
with special cases ${Q}:={Q}_{1}$, ${Y}:={Q}_{0.5}$ (Yule,, 1912) and $H:={Q}_{0.75}$ (Digby,, 1983). The generalized Yule’s ${Q}_{g}$ also has a one-to-one relation to the odds ratio and can be shown to be proper along the same lines as Yule’s ${Q}$. Further, $|{Q}_{g}(A,B)|$ is weakly increasing in $g$ and strictly increasing if ${Cov}(A,B)\neq 0$, which can be seen by computing the derivative with respect to $g$.
We discuss a further property of ${Q}$, ${Q}_{g}$ and the odds ratio, which has been interpreted as a form of invariance to the marginal distributions, in Appendix A.3.
@@
5.3 Comparison of $\phi$, ${C}$ and ${Q}$
It is possible to directly compare the discussed dependence measures in terms of the values they take since they only depend on the marginal event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$ and the joint probability $\mathbb{P}(A\cap B)$ or the value of one of the competing dependence measures, respectively. In Figure 2 we plot Yule’s ${Q}$ and the phi coefficient for a fixed value of Cole’s ${C}$ of 0.7 over all combinations of event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$.
Figure 2: This figure depicts Cole’s ${C}$ with a fixed value of 0.7 and the corresponding values of Yule’s ${Q}$ and the phi coefficient for all combinations of marginal event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$.
@@ 5.3 Comparison of $\phi$, ${C}$ and ${Q}$
Secondly, even though the two proper measures are much closer to each other, they of course also take different values in general. In particular, for equal and at the same time very small or large event propabilities, Yule’s ${Q}$ takes values close to 1 instead of close to 0.7.
For different values of Cole’s coefficient qualitatively the same picture arises. We show the cases where it takes the values 0.3 and $-0.7$ in Figures 11 and 12 in the Appendix. The latter figure in comparison with Figure 2 also illustrates the symmetry relation $\delta(A,\overline{B})=-\delta(A,B)$ formulated in the second part of property (E), which all the coefficients possess.

### v3

@@ 3 Desirable Properties of Dependence Measures
Usually, directed dependence measures lie in the interval [1,1][-1,1] (property (A)), where negative values indicate negative and positive values indicate positive dependence. (B) and (C) make sure that $\delta(A,B)$ correctly indicates the extreme cases of independence as well as perfect positive and negative dependence. Note that, contrary to the case of two general random variables, a dependence measure in the binary case is not a summary measure in the sense of having to condense information (a complicated dependence structure characterized fully by a joint CDF or copula) into a single number, but it usually characterizes the full dependence structure in this single number in the sense that given the marginal event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$ and the dependence measure $\delta(A,B)$, we can recover $\mathbb{P}(A\cap B)$ (indeed, this is possible for all measures discussed in this paper except for the contingency coefficients). As a consequence of this, the “only if” in (B) is infeasible for directed measures of dependence for arbitrary random variables (Embrechts et al.,, 2002), but in the binary case it is feasible. (D) is a very natural property, making sure that stronger dependence leads to more extreme values of $\delta(A,B)$. It is related to the property of coherence proposed by Scarsini, (1984). (C) and (D) ensure that a dependence measure is a sensible indicator of strength of dependence in that larger absolute values of the measure indicate stronger dependence and absolute values close to 0 and 1 indicate weak and strong dependence, respectively. Properties (B) and (D) also imply that the measure correctly indicates the direction of dependence as formalized through the following Proposition.
@@
For a dependence measure $\delta(A,B)$ satisfying $(B)$ and $(D)$ from Definition 3.2, it holds that $\delta(A,B)\geq(\leq)\ 0$ if and only if $A$ and $B$ are positively (negatively) dependent.
@@ 4.1 Covariance
They imply $\frac{1}{4}\leq{Cov}(A,B)\leq\frac{1}{4}$. Thus, property (A) is fulfilled, but property (C) is not. Properties (B) and (D) follow by definition of independence and stronger positive (negative) dependence, (E) directly by representation (1).
@@
Cov(A,B){Cov}(A,B) fulfills properties (A), (B), (D) and (E) from Definition 3.2, but not (C).
Thus, if one found a suitable normalization that made the covariance attainable, a proper dependence measure would arise. Indeed, the phi coefficient, Cole’s C{C} and Yule’s Q{Q}, which we discuss below, all turn out to be normalized versions of the covariance.
@@ 4.2 Phi Coefficient
The phi coefficient is just the classical Pearson correlation coefficient (as well as grade correlation and Kendall’s $\tau_{b}$, which all coincide in this setting) of $\mathds{1}_{A}$ and $\mathds{1}_{B}$ and thus a normalized version of Cov(A,B){Cov}(A,B).
@@
As these bounds are in general not equal to 1-1 and 11 and depend on the marginal event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$, the phi coefficient is not attainable.
@@
Figure 1 visualizes the bounds from (3) as a function of the marginal event probabilities. It shows that the values of the phi coefficient under perfect positive and negative dependence can be very far away from 11 and 1-1. This seriously compromises the interpretability of this measure as it cannot reliably indicate strength of dependence, which should be indicated by $|\phi(A,B)|$ being close to 1. For most combinations of $\mathbb{P}(A)$ and $\mathbb{P}(B)$ such values cannot be reached and instead very small absolute values arise under strong dependence. The phi coefficient is only attainable for the special case of $\mathbb{P}(A)=\mathbb{P}(B)=0.5$ (see Proposition 4.5 below). This shortcoming may be surprising given that the phi coefficient is just a special case of the popular and widely-used Pearson correlation. However, Pearson correlation has serious attainability problems itself and should be used with care (Embrechts et al.,, 2002). In the binary case, they become particularly serious (as two binary random variables are particularly far away from fulfilling the conditions for attainability of Pearson correlation formulated in Fissler and Pohle, (2023, Lemma 3.1), that is, from being symmetric and of the same type).
@@
Thus, the phi coefficient is not a particularly useful measure of dependence for events or binary random variables. Instead, it may be useful for a different task, which becomes clear by analyzing under which conditions $\phi$ takes the values 1-1 and 1.
@@
    AA and $B$ are perfectly positively (negatively) dependent and $\mathbb{P}(A)=\mathbb{P}(B)$ ($\mathbb{P}(A)=\mathbb{P}(\overline{B})$).
@@
What has been recognized in the literature (e.g. Bishop et al., (2007)), is that the phi coefficient equals 1(1-1) under condition (iv), that is, if there are two zeros on the primary (secondary) diagonal of the contingency table, while other measures like Yule’s Q{Q} (in fact, all proper measures by (C) from Definition 3.2 and (iii) in Proposition 2.4) require only one zero on the primary (secondary) diagonal to equal 1 (-1). However, the underlying reasons for this as made clear by Propositions 2.4 and 4.5 have not been uncovered.
The items (vi) and (vii) in Proposition 4.5 are interesting because they establish the connection to conditions well-known from the theory of Pearson correlation and rank correlations, of which $\phi$ is a special case. (vii) reproduces the well-known condition that Pearson correlation is 1 (-1) if and only if the two variables are linear functions of each other, (vi) the well-known condition that grade correlation and Kendall’s $\tau_{b}$ are 1 (-1) if and only if one variable is a monotonic function of the other (Embrechts et al.,, 2002). Thus, equality between the events (or one event and the complement of the other) corresponds to positive (negative) monotonic and linear (the two coincide in the binary case) dependence. Note that neither linear nor monotonic (nor arbitrary) functional relations are useful concepts of perfect dependence in the binary case (or more generally in the discrete case) as a functional relation is only possible under the restriction from (iii) on the marginal distributions, which is in general not fulfilled. This is also the reason why Chatterjee’s $\xi$ is an improper measure (see Appendix A.5).
@@ 4.3 Contingency Coefficients
Contingency coefficients aim at measuring dependence inherent in a general $k\times l$ contingency table. For two categorical random variables $X$ and $Y$, where $X$ takes values $x_{1},...,x_{k}$ and $Y$ takes values $y_{1},...,y_{l}$, Pearson’s Mean Square Contingency (MSC{MSC}) coefficient (Pearson,, 1904) is just the population analogue of the test statistic of Pearson’s chi-squared test of independence:
@@
Contingency coefficients arise by normalizing MSC(X,Y){MSC}(X,Y). As they are designed for nominal random variables, direction of dependence is not a sensible concept and they map to $[0,1]$, only trying to indicate strength of dependence. However, as discussed above, binary random variables can always be regarded as ordinal. Thus, in the $2\times 2$ case, measures of dependence that indicate strength and direction of dependence can be used and the question arises why one should throw away the information on direction of dependence. Since they are widely-used in the binary case, we nevertheless discuss them. In Appendix A.1, we consider Cramérs V, Tschuprow’s T, and Pearson’s contingency coefficient and show that they are closely related to the phi coefficient in the binary setting (for example, the first two equal $|\phi(A,B)|$) and thus inherit its deficiencies, in particular not being proper dependence measures. The same holds true for distance correlation (which also equals $|\phi(A,B)|$) and Chatterjee’s $\xi$ (which equals $\phi^{2}(A,B)$), dependence measures that have recently gained popularity in the statistical literature, see the discussion in Appendix A.5.
@@ 5.1 Cole’s Coefficient
Cole’s Coefficient was proposed by Cole, (1949). An equivalent reformulation of it arises naturally from our discussion of dependence concepts: From the Fréchet–Hoeffding bounds of $\mathbb{P}(A\cap B)$ from Proposition 2.4 arise the bounds on Cov(A,B){Cov}(A,B) from (2), which are attained only under perfect positive and negative dependence of $A$ and $B$ and whose absolute values are in general different from each other. Thus, it is natural to distinguish the cases of positive and negative dependence and normalize covariance with (the absolute values of) those bounds.
@@
The idea of normalizing with the cases of perfect positive and negative dependence has later been used for rank correlations (Vandenhende and Lambert, (2003), Genest and Nešlehová, (2007)) and for Pearson correlation and generalizations of it by Fissler and Pohle, (2023). Indeed, Cole’s C{C} is a special case of attainable modifications of Kendall’s $\tau$, Spearman’s $\rho$ and Pearson correlation as well as of threshold correlation (see Genest and Nešlehová, (2007) for the former two and Fissler and Pohle, (2023) for the latter two).
@@ 5.2 Yule’s Q, the Odds Ratio and Relatives
Besides the phi coefficient and Cole’s coefficient, Yule’s Q{Q} can be viewed as yet another approach to normalizing covariance. It just replaces the minus in the alternative representation of Cov(A,B){Cov}(A,B) from (1) with a plus and normalizes by the resulting term.
Definition 5.3 (Yule’s Q{Q}).
Yule’s Q{Q} is defined as
@@
Considering (6), it is clear that Q(A,B){Q}(A,B) lies in [1,1][-1,1] and fulfills the symmetry properties (E). Its attainability directly follows by invoking part (iii) of Proposition 2.4. The independence property (B) is inherited from the covariance (since the normalization is nonzero). Monotonicity follows from the strictly monotonic relation between Yule’s Q{Q} and the odds ratio (Lemma 5.6) and the monotoniticy of the odds ratio (part (D) of Proposition A.3).
@@
Q(A,B){Q}(A,B) is proper.
The normalization of Yule’s Q{Q} is rather ad hoc, in particular compared to the natural normalization of Cole’s C{C}. The normalization does not use the Fréchet–Hoeffding bounds for covariance, but continuously changes with the joint probabilities. Thus, the measure is not as nicely interpretable in terms of covariance relative to its values under perfect positive or negative dependence. On the other hand, it avoids the need to use a case distinction in the normalization as Cole’s C{C} does, which makes asymptotic theory and inference much easier (see Section 6). One way to interpret Yule’s Q{Q} is by realizing that it is a special case of Goodman-Kruskal’s $\gamma$ (Goodman and Kruskal,, 1954) and can therefore be interpreted in the same vein (Bishop et al.,, 2007).
A very natural way to motivate and interpret Yule’s Q{Q} is via its relation to the odds ratio. The odds ratio is different from the other measures of dependence discussed so far in that it does not build on the covariance and does not fall into the class of classical dependence measures mapping to [1,1][-1,1] or $[0,1]$ as it maps to the extended nonnegative real numbers. Even though it thus falls a bit out of the framework of this paper, we nevertheless discuss it here shortly due to its popularity and to it being nicely interpretable and having nice properties.
@@
We discuss further details on the motivation behind the odds ratio, its interpretation and popularity in Appendix A.2. Note that the odds ratio is closely related to the covariance in that it just replaces the subtraction in (1) by a division. Fittingly, it is also sometimes called the cross-product ratio, referring to the underlying contingency table. The odds ratio and Yule’s Q{Q} are closely related: They are strictly monotonic functions of each other.
@@
Thus, Yule’s Q{Q} can be viewed as a transformation of the odds ratio to a new scale. Hence, it is unsurprising that the odds ratio is proper in an appropriate sense as well. Indeed, we show that it fulfills a modified set of axioms adopted to its codomain and its multiplicative rather than additive nature in Proposition A.3 in the appendix.
Furthermore, Yule’s Q{Q} has been generalized in the sense of
@@
with special cases ${Q}:={Q}_{1}$, ${Y}:={Q}_{0.5}$ (Yule,, 1912) and $H:={Q}_{0.75}$ (Digby,, 1983). The generalized Yule’s ${Q}_{g}$ also has a one-to-one relation to the odds ratio and can be shown to be proper along the same lines as Yule’s Q{Q}. Further, $|{Q}_{g}(A,B)|$ is weakly increasing in $g$ and strictly increasing if ${Cov}(A,B)\neq 0$, which can be seen by computing the derivative with respect to $g$.
We discuss a further property of Q{Q}, ${Q}_{g}$ and the odds ratio, which has been interpreted as a form of invariance to the marginal distributions, in Appendix A.3.
@@
5.3 Comparison of $\phi$, C{C} and Q{Q}
It is possible to directly compare the discussed dependence measures in terms of the values they take since they only depend on the marginal event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$ and the joint probability $\mathbb{P}(A\cap B)$ or the value of one of the competing dependence measures, respectively. In Figure 2 we plot Yule’s Q{Q} and the phi coefficient for a fixed value of Cole’s C{C} of 0.7 over all combinations of event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$.
Figure 2: This figure depicts Cole’s C{C} with a fixed value of 0.7 and the corresponding values of Yule’s Q{Q} and the phi coefficient for all combinations of marginal event probabilities $\mathbb{P}(A)$ and $\mathbb{P}(B)$.
@@ 5.3 Comparison of $\phi$, C{C} and Q{Q}
Secondly, even though the two proper measures are much closer to each other, they of course also take different values in general. In particular, for equal and at the same time very small or large event propabilities, Yule’s Q{Q} takes values close to 1 instead of close to 0.7.
For different values of Cole’s coefficient qualitatively the same picture arises. We show the cases where it takes the values 0.3 and 0.7-0.7 in Figures 11 and 12 in the Appendix. The latter figure in comparison with Figure 2 also illustrates the symmetry relation $\delta(A,\overline{B})=-\delta(A,B)$ formulated in the second part of property (E), which all the coefficients possess.